OpenStudy (zuha):

The ordered pairs(2,-21) and (5,-45) are solutions to which of the following equations: Y=-8x-5 Y=-8x+5 Y=8x-5 Y=8x+5

OpenStudy (jdoe0001):

have you covered slopes yet?

OpenStudy (zuha):

I think so

OpenStudy (p0sitr0n):

plugin x and y values into the equations and see which one gives a true equality

OpenStudy (zuha):

Please guide me

OpenStudy (jdoe0001):

so if those 2 points belong to whatever equation that means that you'd want to find the equation that passes through (2,-21) and (5,-45) thus \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ -21}})\quad &({\color{red}{ 5}}\quad ,&{\color{blue}{ -45}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \) will give you the equation

OpenStudy (zuha):

Ok let me check

OpenStudy (zuha):

P0stir0N@ I'm plugging in those digits into the equation, but It won't come out right

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (2,-21)..x1 = 2 and y1 = -21) (5,-45)..x2 = 5 and y2 = -45 now sub slope(m) = (-45 - (-21) / (5 - 2) slope(m) = (-45 + 21) / 3 slope(m) = -24/3 slope(m) = - 8 with me so far ?

OpenStudy (texaschic101):

let me know and I will continue

OpenStudy (p0sitr0n):

you know the first coordinate is x and the second is y , right?

OpenStudy (texaschic101):

yes...first one is x and second one is y

OpenStudy (zuha):

P0stir0n@ Yes

OpenStudy (zuha):

texaschic101@ sorry I don't understand

OpenStudy (texaschic101):

where are you getting lost

OpenStudy (p0sitr0n):

(2,-21) | | x y

OpenStudy (zuha):

texaschic101 @ I jut completely get lot front he beginning of what you wrote

OpenStudy (zuha):

P0sitr0n @ Yes I know that

OpenStudy (texaschic101):

you are given 2 sets of numbers...set 1 and set 2 label your numbers...set 1 = (x1,y1)....so (2,-21) : x1 = 2 and y1 = -21 set 2 ....(x2,y2) ... (5,-45)...x2 = 5 and y2 = -45 do you understand this

OpenStudy (zuha):

texaschic101 @ yes. i understand

OpenStudy (texaschic101):

now to find the equation, we first need to find the slope. We will use the slope formula for this. slope(m) = (y2 - y1) / (x2 - x1) (2,-21)..x1 = 2 and y1 = -21 (5,-45)..x2 = 5 and y2 = -45 now we sub this into our formula slope(m) = (-45 - (-21) / (5 - 2) slope(m) = (-45 + 21) / 3 slope(m) = -24/3 slope(m) = -8 any questions ? Any at all ?

OpenStudy (texaschic101):

this is not the answer...there is more. I just want to make sure you are understand ing what I am doing. Do you understand ?

OpenStudy (zuha):

I think I understand

OpenStudy (texaschic101):

Now we will continue.. Now you can use the formula jdoe posted up there, but I prefer to just use y = mx + b formula. y = mx + b slope(m) = -8 you can use either of your points...I will use (2,-21)..x = 2 and y = -21 now we sub -21 = -8(2) + b -21 = -16 + b -21 + 16 = b - 5 = b so our slope(m) = -8 and our y intercept (b) = -5....so our equation is : y = -8x - 5 Any questions ?

OpenStudy (zuha):

texaschic101 @ nope

OpenStudy (texaschic101):

ok....let me try again we already found that the slope (m) was -8 now we will use y = mx + b.....we are now looking for b, the y intercept we sub in our slope(m) = -8 we then pick one of our sets of points...(2,-21)...x = 2 and y = -21 now we sub our stuff into the formula. y = mx + b -- solve for b, the y intercept -21 = -8(2) + b -21 = -16 + b -21 + 16 = b - 5 = b now our slope(m) = -8 and our y intercept(b) = -5 so our equation is : y = -8x - 5 Still lost ?

OpenStudy (zuha):

nope, I understand now, thanks

OpenStudy (texaschic101):

good to hear :)